# Identifying Crystal Structures Beyond Known Prototypes from X-ray Powder Diffraction Spectra

### v3

@@ II.1 Wyckoff prototypes and protostructures
We furthermore define that these IDs should be normalized over all possible AFLOW labels that represent the same prototype via transformations of the positions under the coset representatives of the affine normalizer of the spacegroup [105] to use the transformed set whose ID would appear first in lexicographic order. The normalization of our IDs turns them into an origin-independent representation of the respective prototype and protostructure.
@@ II.2 Matching simulated and experimental XRD
Structure matching with experimental XRD data is done by identifying a suitable crystal structure and computing a simulated XRD of this reference structure. Then, the R-value is optimized by perturbing the atoms in the reference structure R-value [85, 86, 96]. In this work, we adopt the approach used by Ward et al. [96, 7] to calculate the R-value, which is given by
@@ IV.1 Generating candidate protostructures
AflowStringGenerator takes chemical composition and spacegroup and returns a list of Wyckoff protostructures. This decoration is constructed by formulating the combinatorial search as a modified subset sum problem. A recursive search function is implemented, which iterates over the multiplicities of Wyckoff sites for a given spacegroup and checks if any combination of summed multiplicity values would be equal to the desired composition (number of atoms). More details on the algorithm for the enumeration are given in the supplementary materials, Ref. 114.
@@ VII Tests of synthetic XRD inversion
To evaluate the performance of the complete workflow in identifying the atomic positions belonging to a previously unseen prototype from an XRD pattern, we randomly chose a few materials from the materials project dataset. These materials were part of the $AB_{2}C_{6}$ space that is excluded from the training data in the retrained models discussed in the previous section. We chose structures that are evaluated to be on the convex hull of thermodynamical stability. Then, we simulated the XRD patterns for these structures. We were able to reproduce the original crystal structure from which the XRD was generated by starting only from the simulated XRD and the composition. The two examples are discussed in more detail below, with a summary presented in Table 1 (more data is available in the supplementary materials, Ref. 114).
@@
Our tests have been run on an NVIDIA A40 GPU, and the Wren formation energy predictions are found to generally, on average, take approximately 0.002 seconds per protostructure. The processing with StructureSolver takes around 30 seconds per protostructure. The computational load can be scaled up more or less perfectly in parallel onto multiple GPUs. This is helpful for potentially solving a large number of XRD peaks that are indexed to sufficient accuracy. Analysis of an additional 25 structures is provided in the supplementary materials, Ref. 114.
@@ VIII Tests of experimental XRD inversion
Figure 4: A few examples of structures identified from experimental XRD patterns in the RRUFF dataset. Each row corresponds to an entry in RRUFF dataset, showing (first column) all the candidate structures identified by our algorithm; (second column) the crystal structure in the RRUFF dataset, regarded here as the true one; (third colmn) comparison of the synthetic XRD from our solutions and the experimental XRD in RRUFF. For the fourth row, our algorithm found two solutions with the same good R-value of 0.016. For the fifth row, we found 5 solutions and only show the one with the lowest R-value here (the others are included in the supplementary materials [114]).
@@ VIII.1 Test of Recovery of structures in RRUFF
We start from the RRUFF project database retrieved on Jan 24, 2024, with 2902 XRD patterns. The dataset is filtered to remove disordered structures (1992), structures that are difficult to automatically process (e.g., internal inconsistencies, unusual symmetry settings, or where the RRUFF data indicates low confidence; 447) and structures of complexity beyond what our present implementation can handle (163) (see Sec. V). This leaves 300 entries. We used DFT calculations using the PBE functional [126] to relax all these structures, and kept only the ones we could confirm to have a resonable stability at zero temperature (i.e., less than 100 meV/atom above the Materials Project [9] convex hull of thermodynamical stability). The end result is a data set of 189 structures suitable for testing our algorithm. (More details on the filtering are available in the supplementary materials, Ref. 114).
Our implementation enumerates and identifies crystal structure candidates for the remaining 189 XRD patterns. We count the inversion as a success if the known structure in RRUFF appears anywhere in our final remaining candidates after applying the energy screening cutoff of 40 meV/atom above the energy of the structure with the lowest predicted energy and has an $R$-value match better than a chosen cutoff of less than 0.1. The result is that 186 out of the 189 structures are recovered. Figure 4 shows some examples of these recovered structures. However, we stress that this test only demonstrates the ability of our algorithm to reproduce structures from experimental XRD that are similar to those available in the training data set (which is very likely to hold for all the RRUFF structures). Also, the workflow yields multiple candidate structures for a given XRD pattern, (e.g. Ilmenite_R130214-9_11917 in Fig 4), which required additional DFT evaluations. More details on this test are available in the supplementary materials, Ref. 114.
@@ VIII.2 Recovery of unidentified ICDD structures
In this section, we investigate previously unidentified experimental XRD data from the International Centre for Diffraction Data (ICDD) database. The edition we use omits the precise atomic coordinates for all entries and does not clearly specify which structures the coordinates are known for. Hence, our selection of relevant XRD to investigate has used the patterns marked as unidentified in the work presented by Griesemer et al. [98]. We detail four cases here (summarized in Table 2), and results for additional ICDD XRD patterns are provided in the supplementary materials, Ref. 114.
@@
A unique Wyckoff protostructure ID does not absolutely ensure that there is no other way to reach the same structure through the already known prototypes, for example, via sub and supergroup relationships and by realizing the crystal structure by placing atoms within the degrees of freedom just slightly on- or off-symmetry in a different prototype. Hence, we calculated formation energies with DFT of all the relevant substitutions into alternative candidate Wyckoff prototypes from both the materials project and WBM to ensure that no other structure reproduced a similar or lower formation energy. We conclude that the found prototype is a new prototype that is not present in those databases. However, we find a very similar prototype ABC2D6_oP20_34_a_b_ab_3c (Materials Project entry: mp-8673), with a comparable R-value of 0.102 which from DFT calculations comes out as slightly less stable at 36 meV/atom above the convex hull. A deeper look into the atomic placements reveals a layered nature of these structures, and it appears the two prototypes represent different stackings between Sb and Fe atoms. (These differences are elaborated on in the supplementary materials, Ref. 114, including a radial distribution plot.)
$\mathrm{SnTe}_{2}\mathrm{Ni}_{3}$ (ICDD ID: 00-060-0238): The ICDD entry specifies spacegroup 194 (P63/mmc). Utilizing the AflowStringGenerator method, we identified a total of 74 unique Wyckoff prototypes. The protostructure with ID A3BC2_hP12_194_ae_c_f:Ni-Sn-Te was determined to be the solution. The calculated crystal structure based on this protostructure is 7 meV/atom above the convex hull of thermodynamical stability from the Materials Project, with an R-value of 0.6. None of the prototypes present in these datasets met either the R-value or the formation energy criteria. The entire computational workflow was completed in a span of 17 minutes.
@@
$\mathrm{In}_{2}\mathrm{Se}_{3}$ (ICDD ID: 00-040-1408): the pattern files were provided and were identified to be in spacegroup P63/mmc. Using our algorithm, we identify that there are two formula units per unit cell, and the protostructure with ID A2B3_hP10_194_f_af:In-Se was determined to be the solution, with 2 degrees of freedom along the z direction at the Wyckoff position f, which was occupied by In (f) and Se (a,f). There were only 16 different combinations of Wyckoff positions that satisfied the composition. The R-value is 0.285, with an energy above the hull of 55 meV/atom. The corresponding Wyckoff prototype (A2B3_hP10_194_f_af) is previously reported, as seen in mp-1094034 and mp-1246153.
$\mathrm{TiFeCl}_{3}$ (ICDD ID: 00-024-1293). The structure was reported to be in spacegroup 186 (P63mc). Upon enumeration, there were a total of 1874 possible Wyckoff protostructures, from which we identified A3BC_hP10_186_c_a_b:Cl-Fe-Tl as the correct protostructure label for the structure with an R-value of 0.069, and an energy of 5 meV/atom below the convex the hull. The prototype corresponding to this structure turns out to already exist in the Materials Project (mp-1096852, mp-19241, etc.) and OQMD. Hence, it was also identified by Griesemer et al. [98] with a formation energy insignificantly higher ($3\ \mathrm{meV}/\mathrm{atom}$) than our structure. (There is a slight complication here in that while they also report their match to be in spacegroup P63mc, their final atomic coordinates differ slightly from ours in a way that makes us identify their structure to be in a supergroup, spacegroup 194 (P63/mmc). Hence, it is possible that the structure actually is in this supergroup rather than the spacegroup reported in ICDD.) A comparison of simulated and experimental patterns for these identifcations is shown in Fig. 3.
@@ IX Discussion
The ability in tests on synthetic patterns to identify prototypes in the structural space omitted in training, along with the identification from previously unidentfied experimental XRD patterns from ICDD of two crystal structures with Wyckoff prototypes not present in existing databases, substantiates the capability of our method to identify new, previously unseen, prototypes. The success in applying the method to experimental patterns also demonstrates robustness against noise and other imperfections absent in simulated patterns. Therefore, our approach shows promise not only in theoretical simulations but also in its applicability to real-world experimental data.

### v4

@@ II.1 Wyckoff prototypes and protostructures
We furthermore define that these IDs should be normalized over all possible AFLOW labels that represent the same prototype via transformations of the positions under the coset representatives of the affine normalizer of the spacegroup [105] to use the transformed set whose ID would appear first when sorted on the sum of the alphabetical indices of the Wyckoff letters and, second, lexicographically on the order they appear in the ID. The normalization of our IDs turns them into an origin-independent representation of the respective prototype and protostructure.
@@ II.2 Matching simulated and experimental XRD
Structure matching with experimental XRD data is done by identifying a suitable crystal structure and computing a simulated XRD of this reference structure. Then, the value of an R-factor is optimized by perturbing the atoms in the reference structure [85, 86, 96]. The $R$-factor is a unitless measure of a relative mismatch between the peaks of the simulated and experimental patterns, but in the context of refining structures, multiple definitions are in use. To allow direct comparisons with Ref. [98] we use their definition:
@@
The context largely determines what can be considered a good or bad $R$-value. Comparing $R$-values across different studies is challenging due to differences in the definitions. Previous works that discuss this topic more broadly suggest that a low $R$-value does not necessarily guarantee a correct structure, nor does a high $R$-value disqualify a candidate [109, 110, 98]. Therefore, it is recommended to visually inspect how well the patterns match [111, 112]. In this work, we primarily use the $R$-value as a guide for quick relative comparisons. Our optimization workflow employs a continuous cost function that is inspired by the $R$-factor (see Sec. III.2).
@@ IV.1 Generating candidate protostructures
AflowStringGenerator takes chemical composition and spacegroup and returns a list of Wyckoff protostructures. This decoration is constructed by formulating the combinatorial search as a modified subset sum problem. A recursive search function is implemented, which iterates over the multiplicities of Wyckoff sites for a given spacegroup and checks if any combination of summed multiplicity values would be equal to the desired composition (number of atoms). More details on the algorithm for the enumeration are given in the supplementary materials, Ref. 118.
@@ VII Tests of synthetic XRD inversion
To evaluate the performance of the complete workflow in identifying the atomic positions belonging to a previously unseen prototype from an XRD pattern, we randomly chose a few materials from the materials project dataset. These materials were part of the $AB_{2}C_{6}$ space that is excluded from the training data in the retrained models discussed in the previous section. We chose structures that are evaluated to be on the convex hull of thermodynamical stability. Then, we simulated the XRD patterns for these structures. We were able to reproduce the original crystal structure from which the XRD was generated by starting only from the simulated XRD and the composition. The two examples are discussed in more detail below, with a summary presented in Table 1 (more data is available in the supplementary materials, Ref. 118).
@@
Our tests have been run on an NVIDIA A40 GPU, and the Wren formation energy predictions are found to generally, on average, take approximately 0.002 seconds per protostructure. The processing with StructureSolver takes around 30 seconds per protostructure. The computational load can be scaled up more or less perfectly in parallel onto multiple GPUs. This is helpful for potentially solving a large number of XRD peaks that are indexed to sufficient accuracy. Analysis of an additional 25 structures is provided in the supplementary materials, Ref. 118.
@@ VIII Tests of experimental XRD inversion
Figure 4: A few examples of structures identified from experimental XRD patterns in the RRUFF dataset. Each row corresponds to an entry in RRUFF dataset, showing (first column) all the candidate structures identified by our algorithm; (second column) the crystal structure in the RRUFF dataset, regarded here as the true one; (third column) comparison of the synthetic XRD from our solutions and the experimental XRD in RRUFF. For the fourth row, our algorithm found two solutions with the same good R-value of 0.012. For the fifth row, we found 5 solutions and only show the one with the lowest R-value here (the others are included in the supplementary materials [118]).
@@ VIII.1 Test of Recovery of structures in RRUFF
We start from the RRUFF project database retrieved on Jan 24, 2024, with 2902 XRD patterns. The dataset is filtered to remove disordered structures (1992), structures that are difficult to automatically process (e.g., internal inconsistencies, unusual symmetry settings, or where the RRUFF data indicates low confidence; 447) and structures of complexity beyond what our present implementation can handle (163) (see Sec. V). This leaves 300 entries. We used DFT calculations using the PBE functional [130] to relax all these structures and kept only the ones we could confirm to have reasonable stability at zero temperature (i.e., less than 100 meV/atom above the Materials Project [9] convex hull of thermodynamical stability). The end result is a data set of 189 structures suitable for testing our algorithm. (More details on the filtering are available in the supplementary materials, Ref. 118).
Our implementation enumerates and identifies crystal structure candidates for the remaining 189 XRD patterns. We count the inversion as a success if the known structure in RRUFF appears anywhere in our final remaining candidates after applying the energy screening cutoff of 40 meV/atom above the energy of the structure with the lowest predicted energy and has an $R$-value match better than a chosen cutoff of less than 0.1. The result is that 186 out of the 189 structures are recovered. Figure 4 shows some examples of these recovered structures. However, we stress that this test only demonstrates the ability of our algorithm to reproduce structures from experimental XRD that are similar to those available in the training data set (which is very likely to hold for all the RRUFF structures). Also, the workflow yields multiple candidate structures for a given XRD pattern, (e.g. Ilmenite_R130214-9_11917 in Fig 4), which required additional DFT evaluations. More details on this test are available in the supplementary materials, Ref. 118.
@@ VIII.2 Recovery of unidentified ICDD structures
In this section, we investigate previously unidentified experimental XRD data from the International Centre for Diffraction Data (ICDD) database. The edition we use omits the precise atomic coordinates for all entries and does not clearly specify which structures the coordinates are known for. Hence, our selection of relevant XRD to investigate has used the patterns marked as unidentified in the work presented by Griesemer et al. [98]. We detail four cases here (summarized in Table 2), and results for additional ICDD XRD patterns are provided in the supplementary materials, Ref. 118.
@@
A unique Wyckoff protostructure ID does not absolutely ensure that there is no other way to reach the same structure through the already known prototypes, for example, via sub and supergroup relationships and by realizing the crystal structure by placing atoms within the degrees of freedom just slightly on- or off-symmetry in a different prototype. Hence, we calculated formation energies with DFT of all the relevant substitutions into alternative candidate Wyckoff prototypes from both the materials project and WBM to ensure that no other structure reproduced a similar or lower formation energy. We conclude that the found prototype is a new prototype that is not present in those databases. However, we find a very similar prototype ABC2D6_oP20_34_a_b_ab_3c (Materials Project entry: mp-8673), with a comparable R-value of 0.102 which from DFT calculations comes out as slightly less stable at 36 meV/atom above the convex hull. A deeper look into the atomic placements reveals a layered nature of these structures, and it appears the two prototypes represent different stackings between Sb and Fe atoms. (These differences are elaborated on in the supplementary materials, Ref. 118, including a radial distribution plot.)
$\mathrm{SnTe}_{2}\mathrm{Ni}_{3}$ (ICDD ID: 00-060-0238): The ICDD entry specifies spacegroup 194 (P63/mmc). Utilizing the AflowStringGenerator method, we identified a total of 74 unique Wyckoff prototypes. The protostructure with ID A3BC2_hP12_194_ae_c_f:Ni-Sn-Te was determined to be the solution. The calculated crystal structure based on this protostructure is 7 meV/atom above the convex hull of thermodynamical stability from the Materials Project, with an R-value of 0.059. None of the prototypes present in these datasets met either the R-value or the formation energy criteria. The entire computational workflow was completed in a span of 17 minutes.
@@
$\mathrm{In}_{2}\mathrm{Se}_{3}$ (ICDD ID: 00-040-1408): the pattern files were provided and were identified to be in spacegroup P63/mmc. Using our algorithm, we identify that there are two formula units per unit cell, and the protostructure with ID A2B3_hP10_194_f_af:In-Se was determined to be the solution, with 2 degrees of freedom along the z direction at the Wyckoff position f, which was occupied by In (f) and Se (a,f). There were only 16 different combinations of Wyckoff positions that satisfied the composition. The R-value is 0.23, with an energy above the hull of 55 meV/atom. The corresponding Wyckoff prototype (A2B3_hP10_194_f_af) is previously reported, as seen in mp-1094034 and mp-1246153.
$\mathrm{TiFeCl}_{3}$ (ICDD ID: 00-024-1293). The structure was reported to be in spacegroup 186 (P63mc). Upon enumeration, there were a total of 1874 possible Wyckoff protostructures, from which we identified A3BC_hP10_186_c_a_b:Cl-Fe-Tl as the correct protostructure label for the structure with an R-value of 0.065, and an energy of 5 meV/atom below the convex the hull. The prototype corresponding to this structure turns out to already exist in the Materials Project (mp-1096852, mp-19241, etc.) and OQMD. Hence, it was also identified by Griesemer et al. [98] with a formation energy insignificantly higher ($3\ \mathrm{meV}/\mathrm{atom}$) than our structure. (There is a slight complication here in that while they also report their match to be in spacegroup P63mc, their final atomic coordinates differ slightly from ours in a way that makes us identify their structure to be in a supergroup, spacegroup 194 (P63/mmc). Hence, it is possible that the structure actually is in this supergroup rather than the spacegroup reported in ICDD.) A comparison of simulated and experimental patterns for these identifcations is shown in Fig. 3.
@@ IX Discussion
The ability in tests on synthetic patterns to identify prototypes in the structural space omitted in training, along with the identification from previously unidentified experimental XRD patterns from ICDD of two crystal structures with Wyckoff prototypes not present in existing databases, substantiates the capability of our method to identify new, previously unseen, prototypes. The success in applying the method to experimental patterns also demonstrates robustness against noise and other imperfections absent in simulated patterns. Therefore, our approach shows promise not only in theoretical simulations but also in its applicability to real-world experimental data.